OpenStudy (anonymous):

How do you solve sin^2x-cos^2x?

OpenStudy (solomonzelman):

It is a double angle identity: \(\large\color{slate}{ \cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b) }\) what happens when \(\large\color{slate}{ a }\) and \(\large\color{slate}{ b }\) are same?

OpenStudy (bossimbacon):

can you help me solomon

OpenStudy (solomonzelman):

you could have messaged, but don't do this in this question, it is his/her question...

OpenStudy (anonymous):

When a and b are the same, it is 1.

OpenStudy (anonymous):

Is that correct?

OpenStudy (solomonzelman):

it is not 1, tell me: \(\large\color{slate}{ \cos(x)\times\cos(x)=? }\) \(\large\color{slate}{ \sin(x)\times\sin(c)=? }\)

OpenStudy (solomonzelman):

For the second one I meant: \(\large\color{slate}{ \sin(x)\times\sin(x) }\)

OpenStudy (anonymous):

cos^2x sin^2x

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

\(\large\color{slate}{ \cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b) }\) when a and b are same, \(\large\color{slate}{ \cos(2x)=\cos(x+x)=\cos(x)\cos(x)-\sin(x)\sin(x) \\=\cos^2(x)\sin^2(x) }\)

OpenStudy (solomonzelman):

oh, the last step another typo

OpenStudy (solomonzelman):

it should say cos^2(x)-sin^2(x)

OpenStudy (solomonzelman):

\(\large\color{slate}{ \cos(2x)=\cos^2(x)-\sin^2(x) }\)

OpenStudy (solomonzelman):

so your answer is?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1-cos^2x